On August 1, Cairle Company’s work-in-process inventory consisted of three jobs with the following costs:

Job 70 Job 71 Job 72
Direct materials $1,700 $2,000 $850
Direct labor 1,900 1,200 900
Applied overhead 1,330 840 630

During August, four more jobs were started. Information on costs added to the seven jobs during the month is as follows:

Job 70 Job 71 Job 72 Job 73 Job 74 Job 75 Job 76
Direct materials $800 $1,235 $3,600 $5,000 $300 $560 $80
Direct labor 1,000 1,400 2,200 1,800 600 900 180

Before the end of August, Jobs 70, 72, 73, and 75 were completed. On August 31, Jobs 72 and 75 were sold.

Required:

1. Calculate the predetermined overhead rate based on direct labor cost.
% of direct labor cost.

2. Calculate the ending balance for each job as of August 31.

Ending Balance
Job 70 $
Job 71 $
Job 72 $
Job 73 $
Job 74 $
Job 75 $
Job 76 $

3. Calculate the ending balance of Work in Process as of August 31.
$

4. Calculate the cost of goods sold for August.
$

5. Assuming that Cairle prices its jobs at cost plus 30 percent, calculate Cairle’s sales revenue for August.
$
Answer #1

1) Predetermine overhead rate = 1330/1900 = 70% of labor cost

2) Job Cost Sheet

Job 70 Job 71 Job 72 Job 73 Job 74 Job 75 Job 76
Beginning balance 4930 4040 2380
Direct material 800 1235 3600 5000 300 560 80
Direct labor 1000 1400 2200 1800 600 900 180
Overhead 700 980 1540 1260 420 630 126
Total 7430 7655 9720 8060 1320 2090 386

3) Ending WIP = 7655+1320+386 = 9361

4) Cost of goods sold = 9720+2090 = 11810

5) Sales revenue = 11810*1.3 = $15353